The resistance of a wire to the transmission of electricity varies directly as the length of the wire and inversely as the square of the radius of a cross section of the wire. An 800m has a radius of 0.48cm and ha a resistance of 0.3125 ohms. if another wire made from the ame metal has a length of 500 m and a resistance of 0.5 ohms, what is the radius?
Answer by stanbon(75887)

The resistance of a wire to the transmission of electricity varies directly as the length of the wire and inversely as the square of the radius of a cross section of the wire.
----
r = kL/r^2
---
Solve for "k" using "An 800m has a radius of 0.48cm and ha a resistance of 0.3125 ohms".
---
0.3125 = k*800/(0.48)^2
---
k = 0.00009
=================
Equation:
r = 0.00009L/r^2
----
If another wire made from the same metal has a length of 500 m and a resistance of 0.5 ohms, what is the radius?
---
r = 0.00009*500/(0.5)^2
---
r = 0.18 ohms
===================
Cheers,
Stan H.